# Eco-evolutionary dynamics of cooperative antimicrobial resistance in a population of fluctuating volume and size

### v1

@@ 3 Static environment
To gain an insight into the model’s eco-evolutionary dynamics, it is instructive to analyse its properties in a static environment. In this section, we thus consider that there is no EV, with the carrying capacity kept constant, $K(t)=K_{0}$.
@@ 3.1 Mean-field: The absence of fluctuations promotes coexistence
It is useful to start with the case of a very large population and carrying capacity $K(t)=K_{0}\gg 1$. In this mean-field setting, we ignore demographic and environmental fluctuations, and the population’s dynamics is aptly described by the mean-field (deterministic) differential equations [WFM17, WFM18]
@@
where the dot indicates the time derivative. It is clear from (7) and (8) that the dynamics of $N$ and $x$ are decoupled at the mean-field level. This stems from having a constant concentration cooperation threshold $x_{\text{th}}$. According to the logistic equation (7), the population size reaches its equilibrium, $N=K_{0}$, on a time scale $t\sim\mathcal{O}(1)$, while equation (8) implies that the population composition has a stable equilibrium $x=x_{\text{th}}$ that is reached on a time scale $t\sim\mathcal{O}(1/s)$ when $x\geq x_{\text{th}}$ and $t\sim\mathcal{O}(1/(a-s))$ when $x<x_{\text{th}}$. Moreover, when $s<a\ll 1$, the metabolic cost $s$ and growth hindrance $a$ are small and there is a timescale separation: $N$ relaxes to $K_{0}$ much faster than $x$ reaches $x_{\text{th}}$. We note that the coexistence equilibrium in terms of $R$ and $S$ is $N_{R}=x_{\text{th}}K_{0}$ and $N_{S}=(1-x_{\text{th}})K_{0}$.
@@ 3.2 The fate of a large population in a static environment is determined by the cooperation threshold
We now consider that the population size is large but finite, and the environment is static, with a constant carrying capacity $K(t)=K_{0}$. In this setting, the long-time dynamics is characterised by the eventual fixation of one of the strains and the extinction of the other.
In a large population, the community composition approaches the coexistence equilibrium, which coincides with the cooperation concentration threshold, i.e., $x\rightarrow x_{\text{th}}$, see equation (8), and the total population reaches its steady state $N\to K_{0}$ before any fixation/extinction events likely occur, see equation (7). We can therefore analyse the fate of a large finite population in a static environment by assuming a constant size coinciding with the carrying capacity, $N=K_{0}$, and that any fixation/extinction events occur from the population coexistence equilibrium $x_{\text{th}}$. This neglects transients of $N$ and $x$ towards $K_{0}$ and $x_{\text{th}}$, respectively.
When the fixed total population $N\approx K_{0}$ is neither too small nor too large ($1\ll K_{0}\sim K_{-}\ll K_{+}$ in our examples), fluctuations are of order $\sqrt{K_{0}}$ and can thus prevent long-lived coexistence by leading to the fast fixation of one strain. Here, we are interested in characterising these scenarios of dominance by one strain or coexistence. For this, the evolutionary dynamics is modelled in terms of the analytically tractable Moran process [Mor62, Ewe04, BM07, CMF11, WFM17, WFM18]. In the Moran stochastic dynamics, each birth/death of one $R$ is balanced by the simultaneous death/birth of one $S$. This ensures that the overall population size in the Moran model remains constant at each update of its composition, as $N=K_{0}$. The simultaneous Moran birth/death events thus occur according to the reactions
@@
with the effective transition rates $\widetilde{T}^{\pm}_{R}\equiv T^{\pm}_{R}T^{\mp}_{S}/N=T^{\pm}_{R}T^{\mp}_{S}/K_{0}$ obtained from (3[WFM17, WFM18]:
@@
$\widetilde{T}^{+}_{R}$ $=\frac{(1-s)x(1-x)}{1-a\theta\left[x_{\text{th}}-x\right]+(a\theta\left[x_{\text{th}}-x\right]-s)x}\frac{N^{2}}{K_{0}},\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and}$
$\widetilde{T}^{-}_{R}$ $=\frac{(1-a\theta\left[x_{\text{th}}-x\right])x(1-x)}{1-a\theta\left[x_{\text{th}}-x\right]+(a\theta\left[x_{\text{th}}-x\right]-s)x}\frac{N^{2}}{K_{0}},$ (9)
where $x\equiv N_{R}/N$ and here $N=K_{0}$ (constant population size), which simplifies the expressions of $\widetilde{T}^{\pm}_{R}$. However, for the discussion of the effect of EV in section 4, it is convenient to keep an explicit $N$ dependence.
@@
Fixation probability. We are interested in the probability that a strain takes over the entire population starting from a given initial number of microbes, a quantity referred to as fixation probability [Ewe04, CK09, AS06, BM07]. Here, the fixation probability of the strain $R$ when the population initially consists of $N_{R}^{0}$ resistant individuals (and $N_{S}^{0}=N-N_{R}^{0}$ sensitive cells) is denoted by $\phi_{N}\left(N_{R}^{0}\right)$, where the subscript signifies that the expression is for a population of constant size $N$. Since the fixation of $R$ coincides with the extinction of $S$, $\phi_{N}\left(N_{R}^{0}\right)$ also gives the extinction probability of the strain $S$.
For the Moran model and a population of constant size $N$, the exact expression of the $R$ fixation probability is [Gar02, vK92, Ewe04, AS06, TH09]
$\phi_{N}\left(N_{R}^{0}\right)=\frac{1+\sum_{k=1}^{N_{R}^{0}-1}\prod_{i=1}^{k}\gamma\left(i,N\right)}{1+\sum_{k=1}^{N-1}\prod_{i=1}^{k}\gamma\left(i,N\right)},\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ \gamma\left(N_{R},N\right)\equiv\frac{\widetilde{T}^{-}_{R}\left(N_{R},N\right)}{\widetilde{T}^{+}_{R}\left(N_{R},N\right)}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ N_{R}^{0}=1,2,...,N,$ (10)
@@
and the quantity $\gamma\left(N_{R},N\right)$ fully determines $\phi_{N}\left(N_{R}^{0}\right)$. When $N$ is sufficiently large, $x_{0}\equiv N_{R}^{0}/N$ is a continuous variable. The fixation probability can thus be regarded as a continuous function of $x_{0}$, and it is convenient to write $\phi_{N}\left(N_{R}^{0}=Nx_{0}\right)$ as $\phi_{N}\left(x_{0}\right)$. According to equation (8), the make-up of a large population in a static environment rapidly reaches its coexistence equilibrium for any initial condition $x_{0}$ well separated from the boundaries, i.e., $x\rightarrow x_{\text{th}}$, and lingers about $x_{\text{th}}$ until fixation/extinction occurs. When $x_{0}$ is not too close to 00 or $1$, we can thus assume $\phi(x_{0})\simeq\phi_{N}\left(x_{0}=x_{\text{th}}\right)$, yielding the following expression for the fixation probability of $R$ in a population of constant size $N$ [HNARM23a]:
$\phi_{N}(x_{\text{th}})=\frac{1-\left(\frac{1-a}{1-s}\right)^{x_{\text{th}}N}}{1-\left(\frac{1-a}{1-s}\right)^{x_{\text{th}}N}+\frac{a-s}{s(1-a)}\left(\frac{1-a}{1-s}\right)^{x_{\text{th}}N}\left[\left(\frac{1}{1-s}\right)^{(1-x_{\text{th}})N}-1\right]}\simeq\frac{1}{1+\left(\frac{1}{1-a}\right)^{-N\left(x_{\text{th}}-x_{\text{th}}^{*}\right)}},$ (11)
@@
with $\text{with\leavevmode\nobreak\ }x_{\text{th}}^{*}\equiv\frac{\ln\left(1-s\right)}{\ln\left(1-a\right)}+\frac{1}{N}\frac{\ln\left(\frac{s(1-a)}{a-s}\right)}{\ln\left(1-a\right)}.$
@@
The compact (approximation) of $\phi_{N}\left(x_{\text{th}}\right)$ in the right-hand-side has been obtained by neglecting $(1-a)^{x_{\text{th}}N}/(1-s)^{x_{\text{th}}N}\ll 1$ and $(1-s)^{(1-x_{\text{th}})N}\ll 1$. Here $x_{\text{th}}^{*}$ is defined as the critical value of the cooperation concentration threshold for which both strains have probability $1/2$ of fixation, i.e. $\phi_{N}\left(x^{*}_{\text{th}}\right)=1/2$. We have found that, for biologically plausible values of $0<s<a\lesssim 10^{-1}$ and $N>25$, the right hand side of equation (11) is a good approximation for the fixation probability and the simplified compact expression matches up well with the exact solution and simulation data; see figure 2(a). Remarkably, equation (11) reveals that strain $R$ is the most likely to go extinct (and $S$ to fixate) when the cooperation threshold is below the critical value $x_{\text{th}}^{*}$: $\phi_{N}(x_{\text{th}})\ll 1$ if $x_{\text{th}}<x_{\text{th}}^{*}$. On the other hand, when the cooperation threshold exceeds $x_{\text{th}}^{*}$, a large concentration of $R$ is thus needed to inactivate the drug, and $R$ is likely to fixate and $S$ to go extinct, since equation (11) predicts $\phi_{N}(x_{\text{th}})\approx 1$ if $x_{\text{th}}>x_{\text{th}}^{*}$, as confirmed by simulations in figure 2(a). As an intriguing feature of the cooperative resistance, we notice that the more efficient in producing resistance enzymes (public good) $R$ is, the lower $x_{\text{th}}$ is, and the more likely $R$ is to go extinct. Similarly, the less efficient in producing the public good $R$ is, the more likely it is to fixate (high $x_{\text{th}}$).
@@
Mean coexistence time. Another quantity of great interest is the mean time until one of the strain fixates and the other is wiped out, referred to as the (unconditional) mean fixation time or, equivalently here, as the (unconditional) mean extinction time. Since the strains here coexist until fixation/extinction of one of them, the (unconditional) mean fixation time here coincides with the mean coexistence time (MCT) of the strains. For the Moran model in a fixed population of fixed size $N$, the exact expression of the MCT is well-known [Gar02, vK92, Ewe04, AS06, TH09], and can here be written as
$\langle\tau_{N}\left(N^{0}_{R}\right)\rangle=\phi_{N}(N^{0}_{R})\sum_{k=N_{R}^{0}}^{N-1}\sum_{n=1}^{k}\frac{\prod_{m=n+1}^{k}\gamma\left(m/N\right)}{\widetilde{T}^{+}_{R}\left(n/N\right)}-\left[1-\phi_{N}(N^{0}_{R})\right]\sum_{k=1}^{N_{R}^{0}-1}\sum_{n=1}^{k}\frac{\prod_{m=n+1}^{k}\gamma\left(m/N\right)}{\widetilde{T}^{+}_{R}\left(n/N\right)},$ (12)
where the subscript indicates that the population size $N$ is fixed (here $N=K_{0}$) and $\phi_{N}(N^{0}_{R})$ is the $R$ fixation probability given by equation (10). As above, since fixation/extinction, generally arises from the coexistence equilibrium $x_{\text{th}}$, we can assume $\langle\tau_{N}\left(x_{0}\right)\rangle\simeq\langle\tau_{N}\left(x_{0}=x_{\text{th}}\right)\rangle$ when $x_{0}$ is not too close to 00 or $1$, and obtain the simplified approximate expression
$\langle\tau_{N}(x_{\text{th}})\rangle\simeq\frac{a(1-s)}{s^{2}(a-s)}\frac{\phi(x_{\text{th}},N)}{x_{\text{th}}}\frac{\left(\frac{1}{1-s}\right)^{(1-x_{\text{th}})N}-1}{(1-x_{\text{th}})N},$ (13)
@@
whose derivation is given in section B of the appendix. This expression of the MCT closely reproduces the exact prediction (12) and simulation data of figure 2(b) for $\langle\tau_{N}\rangle>50$. We notice that $\langle\tau_{N}(x_{\text{th}})\rangle\ll\langle\tau_{N}(x_{\text{th}}^{*})\rangle$ when $x_{\text{th}}$ is much below or above $x_{\text{th}}^{*}$. This means that, for given $s,a$ and $N$, the MCT is maximum for $x_{\text{th}}\simeq x_{\text{th}}^{*}$, and it decreases exponentially as $x_{\text{th}}$ deviates from $x^{*}_{\text{th}}$. This can be interpreted as follows: strongly and weakly cooperative $R$ (high/low $x_{\text{th}}$, respectively) favour the rapid dominance of one strain, whereas $R$ with a cooperative threshold about $x^{*}_{\text{th}}$ promotes long-lived coexistence, and in this case the MCT greatly exceeds the population size [AS06, AHNRM23, AM10, AM11, CRF09, HMT11, RMF07].
@@
Coexistence probability. In section 3.1, we have seen that the mean-field dynamics of an infinitely large population leads to a stable coexistence of the strains. However, coexistence cannot be stable in a finite population since demographic fluctuations unavoidably cause the fixation of one stain and the extinction of the other [MA10, AM10, AHNRM23]. In fact, while we are guaranteed that only one of the strains will finally survive, its fixation can occur after a very long-lived coexistence of the strains. When this happens, the mean-field coexistence equilibrium becomes metastable. The population is thus at quasi-stationarity, and the MCT generally scales superlinearly with the population size $N$, with a fixation/extinction time, here denoted by $\tau_{N}$, that is generally exponentially distributed with a cumulative distribution $1-{\rm exp}(-\tau_{N}/\langle\tau_{N}\rangle)$, see e.g. [MA10, AM10, AM11, AM17, AHNRM23]. This is in stark contrast with the MCT scaling sublinearly with $N$ when one species dominates over the other [AS06, AHNRM23, AM10, AM11, CRF09, HMT11, RMF07]. The scenarios of dominance and long-lived coexistence are hence separated by a regime where the MCT scales linearly with the population size [CRF09, HMT11, RMF07, HNARM23a, AHNRM23]. This leads us to consider that there is long-lived coexistence of $R$ and $S$ strains whenever the coexistence time exceeds $2N$ (where the conservative factor $2$ is chosen for convenience as in [HNARM23a, AHNRM23]), and otherwise one of the strains dominates and fixates in a time $\tau_{N}$ less than $2N$. Here, we will therefore consider that the probability to have long-lived coexistence in a population of size $N$ is $\eta_{N}={\rm Prob.}\left(\tau_{N}>2N\right)$, the probability that coexistence time exceeds $2N$, which is obtained from its cumulative distribution:
@@
The quantities (14) and (15) can be used to obtain fixation-coexistence diagrams and visualise the long-time eco-evolutionary properties of the model in static environments [HNARM23a, AHNRM23]. Based on these results, and assuming ${\rm min}(x_{\text{th}},1-x_{\text{th}})\lesssim 1/\sqrt{N}$ so that demographic fluctuations are strong enough to cause fast fixation/extinction events, the long-time behaviour of a finite population of size $N$ in a static environment (with $sN\gg 1$ and $(a-s)N\gg 1$) can be summarised in terms of the value of $x_{\text{th}}$ relative to $x_{\text{th}}^{*}$. This is, if $x_{\text{th}}<x_{\text{th}}^{*}$, the sensitive strain $S$ dominates, while $R$ most probably fixate when $x_{\text{th}}>x_{\text{th}}^{*}$; whereas long-lived coexistence of the strains is expected when $x_{\text{th}}$ is about $x_{\text{th}}^{*}$. When ${\rm min}(x_{\text{th}},1-x_{\text{th}})\gg 1/\sqrt{N}$, demographic fluctuations are too weak to ensure fast fixation/extinction, and long-lived coexistence is always expected. The above approach and quantities (14) and (15) are generalised to a dynamic environment in section 4.3 to obtain the corresponding fixation-coexistence diagrams, see equations (19) and (24) and figure 4.
Figure 2: $R$ fixation probability and mean coexistence time (MCT) in static environments. (a) $R$ fixation probability $\phi_{N}$ in terms of the concentration cooperation threshold $x_{\text{th}}$ for $s=0.1$, $a=0.25$, and for five examples of total population size, from $N=25$ (yellow green) to 200 (dark green); the starting microbial composition is set at the coexistence equilibrium $x_{0}=x_{\text{th}}$; dotted and dashed lines show the exact and approximated Moran predictions of equation (11), respectively, which are only distinguishable for the smallest population size; noisy solid lines are from simulation data ($10^{3}$ realisations for each data point) and match with dotted lines. The vertical black dashed line shows the value of $x_{\text{th}}^{*}$ for which each strain has a probability $1/2$ to fixate when $N\gg 1$. (b) Mean coexistence time $\langle\tau_{N}\rangle$ vs $x_{\text{th}}$ in log-linear scale; the simplified formula of equation (13) (dashed lines), approximates well (for $\langle\tau_{N}\rangle>50$) the exact MCT (12) (dotted lines) and simulation results for the MCT (solid lines); legend and line styles as in panel (a).

### v2

@@ 3 Static environment
To gain an insight into the model’s eco-evolutionary dynamics, it is instructive to analyse its properties in a static environment. In this section, we thus consider that there is no EV, with the carrying capacity kept constant, $K(t)=K_{0}$. Below we show that the proportion of resistant microbes always tends to the fixed cooperation threshold $x_{\text{th}}$ in static environments. However, when the microbial population is scarce, fluctuations due to birth-death events are significant, and demographic noise can thus be responsible for the extinction or fixation of the resistant strain.
@@ 3.1 Mean-field: The absence of fluctuations promotes coexistence
It is useful to start with the case of a very large population and carrying capacity $K(t)=K_{0}\gg 1$. In this mean-field setting, we ignore demographic and environmental fluctuations, and the population’s dynamics is aptly described by the mean-field (deterministic) differential equations [11, 12]
@@
where the dot indicates the time derivative. It is clear from (7) and (8) that the dynamics of $N$ and $x$ are decoupled at the mean-field level. This stems from having a constant concentration cooperation threshold $x_{\text{th}}$. According to the logistic equation (7), the population size reaches its equilibrium, $N=K_{0}$, on a time scale $t\sim\mathcal{O}(1)$, while equation (8) implies that the population composition has a stable equilibrium $x=x_{\text{th}}$ that is reached on a time scale $t\sim\mathcal{O}(1/s)$ when $x\geq x_{\text{th}}$ and $t\sim\mathcal{O}(1/(a-s))$ when $x<x_{\text{th}}$. Moreover, when $s<a\ll 1$, the metabolic cost $s$ and growth hindrance $a$ are small and there is a timescale separation: $N$ relaxes to $K_{0}$ much faster than $x$ reaches $x_{\text{th}}$. We note that the coexistence equilibrium in terms of $R$ and $S$ is $N_{R}=x_{\text{th}}K_{0}$ and $N_{S}=(1-x_{\text{th}})K_{0}$. This stable coexistence arises from sensitive microbes out-competing resistant ones in the presence of antimicrobial drug, but only if the resistance is shared, which requires a minimum fraction of resistant cells in the population.
@@ 3.2 The fate of a large population in a static environment is determined by the cooperation threshold
We now consider that the population size is large but finite, and the environment is static, with a constant carrying capacity $K(t)=K_{0}$. In this setting, the long-time dynamics is characterised by the eventual fixation of one of the strains due to birth-death demographic fluctuations. Below, we show that for large concentration cooperation thresholds ($10\ll x_{\text{th}}<$ 1) the resistant strain fixates the population quickly, whereas sensitive cells rapidly take over under smaller thresholds ($0<x_{\text{th}}\ll 1$). Moreover, we find the cooperation threshold $x_{\text{th}}^{*}$ at which both strains have the same fixation probability, and obtain a simple expression for the coexistence duration that increases exponentially with the total population size.
As discussed in the previous section, in a large population the community composition rapidly approaches the coexistence equilibrium, which coincides with the cooperation concentration threshold, i.e., $x\rightarrow x_{\text{th}}$, see equation (8), and the total population reaches its steady state $N\to K_{0}$ before any fixation/extinction events likely occur, see equation (7). We can therefore analyse the fate of a large finite population in a static environment by assuming a constant total population size coinciding with the carrying capacity, $N=K_{0}$, and that any fixation/extinction events occur from the population coexistence equilibrium $x_{\text{th}}$.
When the fixed total population $N\approx K_{0}$ is neither too small nor too large ($1\ll K_{0}\sim K_{-}\ll K_{+}$ in our examples), demographic fluctuations are significant (of order $\sqrt{K_{0}}$) and can thus prevent long-lived coexistence by leading to the fast fixation of one strain. Here, we are interested in characterising these scenarios of dominance by one strain or coexistence. For this, the evolutionary dynamics is modelled in terms of the analytically tractable Moran process [75, 33, 35, 77, 11, 12]. In the Moran stochastic dynamics, each birth/death of one $R$ is balanced by the simultaneous death/birth of one $S$. This ensures that the overall population size in the Moran model remains constant at each update of its composition, as $N=K_{0}$. The simultaneous Moran birth/death events thus occur according to the reactions
@@
with the effective transition rates $\widetilde{T}^{\pm}_{R}\equiv T^{\pm}_{R}T^{\mp}_{S}/N=T^{\pm}_{R}T^{\mp}_{S}/K_{0}$ obtained from (3[11, 12]:
@@
$\widetilde{T}^{+}_{R}$ $=\frac{(1-s)x(1-x)}{1-a\theta\left[x_{\text{th}}-x\right]+(a\theta\left[x_{\text{th}}-x\right]-s)x}N,\text{\leavevmode\nobreak\ \leavevmode\nobreak\ and}$
$\widetilde{T}^{-}_{R}$ $=\frac{(1-a\theta\left[x_{\text{th}}-x\right])x(1-x)}{1-a\theta\left[x_{\text{th}}-x\right]+(a\theta\left[x_{\text{th}}-x\right]-s)x}N,$ (9)
where $x\equiv N_{R}/N$ and where we used $N=K_{0}$ (constant population size). Therefore, the stochastic dynamics of $x$, when the total microbial population size is kept constant (static environment), is well described by the above one-dimensional effective birth-death reactions.
@@
Fixation probability. We are interested in the probability that a strain takes over the entire population starting from a given initial number of microbes, a quantity referred to as fixation probability [33, 34, 76, 35]. Here, the fixation probability of the strain $R$ when the population initially consists of $N_{R}^{0}$ resistant individuals (and $N_{S}^{0}=N-N_{R}^{0}$ sensitive cells) is denoted by $\phi_{N}\left(x_{0}\right)$, where $x_{0}\equiv N_{R}^{0}/N$ is the initial fraction of resistant microbes, and where the subscript signifies that the expression is for a population of constant size $N$. Since the fixation of $R$ coincides with the extinction of $S$, $\phi_{N}\left(x_{0}\right)$ also gives the extinction probability of the strain $S$.
@@
For the one-dimensional Moran model (e.g., two microbial strains at a fixed total population $N$), the exact general expression of the fixation probability $\phi_{N}\left(x_{0}\right)$ is well known [72, 73, 33, 76, 74], see equation (S1). According to equation (8), the make-up of a large population in a static environment rapidly reaches its coexistence equilibrium from any moderate initial $R$ fraction ($0\ll x_{0}\ll 1$), i.e., $x_{0}\rightarrow x_{\text{th}}$, and lingers about $x_{\text{th}}$ until fixation/extinction occurs. When $x_{0}$ is not too close to 00 or $1$, we can thus assume $\phi_{N}(x_{0})\simeq\phi_{N}\left(x_{0}=x_{\text{th}}\right)$, yielding the following approximate expression for the fixation probability of $R$ in a population of large constant size $N$ [32]:
@@
$\phi_{N}(x_{\text{th}})\simeq\frac{1}{1+\left(\frac{1}{1-a}\right)^{-N\left(x_{\text{th}}-x_{\text{th}}^{*}\right)}},\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ with\leavevmode\nobreak\ }x_{\text{th}}^{*}\equiv\frac{\ln\left(1-s\right)}{\ln\left(1-a\right)}+\frac{1}{N}\frac{\ln\left(\frac{s(1-a)}{a-s}\right)}{\ln\left(1-a\right)};$ (10)
@@
see appendix section B, where the exact expression of $\phi_{N}(x_{\text{th}})$ is given by equation (S2). Here $x_{\text{th}}^{*}$ is defined as the critical value of the cooperation concentration threshold for which both strains have probability $1/2$ of fixation, i.e. $\phi_{N}\left(x^{*}_{\text{th}}\right)=1/2$. We have found that, for biologically plausible values of $0<s<a\lesssim 10^{-1}$ and $N>25$, equation (10) is a good approximation for the fixation probability and this simplified compact expression matches up well with the exact solution and simulation data; see figure 2(a). Remarkably, equation (10) reveals that strain $R$ is the most likely to go extinct (and $S$ to fixate) when the cooperation threshold is below the critical value $x_{\text{th}}^{*}$: $\phi_{N}(x_{\text{th}})\ll 1$ if $x_{\text{th}}<x_{\text{th}}^{*}$. On the other hand, when the cooperation threshold exceeds $x_{\text{th}}^{*}$, a large concentration of $R$ is thus needed to inactivate the drug, and $R$ is likely to fixate and $S$ to go extinct, since equation (10) predicts $\phi_{N}(x_{\text{th}})\approx 1$ if $x_{\text{th}}>x_{\text{th}}^{*}$. As an intriguing feature of the cooperative resistance, we notice that the more efficient in producing resistance enzymes (public good) $R$ is, the lower $x_{\text{th}}$ is, and the more likely $R$ is to go extinct. Similarly, the less efficient in producing the public good $R$ is, the more likely it is to fixate (high $x_{\text{th}}$).
@@
Mean coexistence time. Another quantity of great interest is the mean time until one of the strain fixates and the other is wiped out. Since the strains here coexist until fixation/extinction of one of them, this (unconditional) mean fixation time coincides with the mean coexistence time (MCT) of the strains $\langle\tau_{N}\left(x_{0}\right)\rangle$. For the Moran model in a population of fixed size $N$, the exact expression of the MCT is well-known [72, 73, 33, 76, 74], see equation (S3). As discussed previously, we can assume that fixation/extinction arises from the coexistence equilibrium $x_{0}=x_{\text{th}}$, and then obtain the simplified approximate MCT expression
$\langle\tau_{N}(x_{\text{th}})\rangle\simeq\frac{a(1-s)}{s^{2}(a-s)}\frac{\phi_{N}(x_{\text{th}})}{x_{\text{th}}}\frac{\left(\frac{1}{1-s}\right)^{(1-x_{\text{th}})N}-1}{(1-x_{\text{th}})N},$ (11)
@@
whose derivation is given in section C of the appendix. This expression of the MCT closely reproduces the exact prediction (S3) and simulation data of figure 2(b) for $\langle\tau_{N}\rangle>50$. We notice that $\langle\tau_{N}(x_{\text{th}})\rangle\ll\langle\tau_{N}(x_{\text{th}}^{*})\rangle$ when $x_{\text{th}}$ is much below or above $x_{\text{th}}^{*}$. This means that, for given $s,a$ and $N$, the MCT is maximum for $x_{\text{th}}\simeq x_{\text{th}}^{*}$, and it decreases exponentially as $x_{\text{th}}$ deviates from $x^{*}_{\text{th}}$. This can be interpreted as follows: strongly and weakly cooperative $R$ (low/high $x_{\text{th}}$, respectively) favour the rapid dominance of one strain, whereas $R$ with a cooperative threshold about $x^{*}_{\text{th}}$ promotes long-lived coexistence, and in this case the MCT greatly exceeds the population size [76, 13, 87, 88, 89, 90, 91].
@@
Coexistence probability. In section 3.1, we have seen that the mean-field dynamics of an infinitely large population leads to a stable coexistence of the strains. However, coexistence cannot be stable in a finite population since demographic fluctuations unavoidably cause the fixation of one stain and the extinction of the other [92, 87, 13]. In fact, while we are guaranteed that only one of the strains will finally survive, its fixation can occur after a very long-lived coexistence of the strains. When this happens, the mean-field coexistence equilibrium becomes metastable. The population is thus at quasi-stationarity, and the MCT generally scales superlinearly with the population size $N$, with a fixation/extinction time, here denoted by $\tau_{N}$, that is generally exponentially distributed with a cumulative distribution $1-{\rm exp}(-\tau_{N}/\langle\tau_{N}\rangle)$, see e.g. [92, 87, 88, 93, 13]. This is in stark contrast with the MCT scaling sublinearly with $N$ when one species dominates over the other [76, 13, 87, 88, 89, 90, 91]. The scenarios of dominance and long-lived coexistence are hence separated by a regime where the MCT scales linearly with the population size [89, 90, 91, 32, 13]. This leads us to consider that there is long-lived coexistence of $R$ and $S$ strains whenever the coexistence time exceeds $2N$ (where the conservative factor $2$ is chosen for convenience as in [32, 13]), and otherwise one of the strains dominates and fixates in a time $\tau_{N}$ less than $2N$. Here, we will therefore consider that the probability to have long-lived coexistence in a population of size $N$ is $\eta_{N}={\rm Prob.}\left(\tau_{N}>2N\right)$, the probability that coexistence time exceeds $2N$, which is obtained from its cumulative distribution:
@@
The quantities (12) and (13) can be used to obtain fixation-coexistence diagrams and visualise the long-time eco-evolutionary properties of the model in static environments [32, 13]. Based on these results, and assuming ${\rm min}(x_{\text{th}},1-x_{\text{th}})\lesssim 1/\sqrt{N}$ (so that demographic fluctuations about the concentration cooperation threshold are strong enough to cause fast fixation/extinction events), the long-time behaviour of a finite population of size $N$ in a static environment (with $sN\gg 1$ and $(a-s)N\gg 1$) can be summarised in terms of the value of $x_{\text{th}}$ relative to $x_{\text{th}}^{*}$. This is, if $x_{\text{th}}<x_{\text{th}}^{*}$, the sensitive strain $S$ dominates, while $R$ most probably fixate when $x_{\text{th}}>x_{\text{th}}^{*}$; whereas long-lived coexistence of the strains is expected when $x_{\text{th}}$ is about $x_{\text{th}}^{*}$. When ${\rm min}(x_{\text{th}},1-x_{\text{th}})\gg 1/\sqrt{N}$, demographic fluctuations are too weak to ensure fast fixation/extinction, and long-lived coexistence is always expected. The above approach and quantities (12) and (13) are generalised to a dynamic environment in section 4.3 to obtain the corresponding fixation-coexistence diagrams, see equations (17) and (22) and figure 4.
Figure 2: $R$ fixation probability and mean coexistence time (MCT) in static environments. (a) $R$ fixation probability $\phi_{N}$ in terms of the concentration cooperation threshold $x_{\text{th}}$ for $s=0.1$, $a=0.25$, and for five examples of total population size, from $N=25$ (yellow green) to 200 (dark green); the starting microbial composition is set at the coexistence equilibrium $x_{0}=x_{\text{th}}$; dotted and dashed lines show the exact and approximated Moran predictions of equations (S2) and (10), respectively, which are only distinguishable for the smallest population size; noisy solid lines are from simulation data ($10^{3}$ realisations for each data point) and match with dotted lines. The vertical black dashed line shows the value of $x_{\text{th}}^{*}$ for which each strain has a probability $1/2$ to fixate when $N\gg 1$. (b) Mean coexistence time $\langle\tau_{N}\rangle$ vs $x_{\text{th}}$ in log-linear scale; the simplified formula of equation (11) (dashed lines), approximates well (for $\langle\tau_{N}\rangle>50$) the exact MCT (S3) (dotted lines) and simulation results for the MCT (solid lines); legend and line styles as in panel (a).